Area of a triangle. Find the exact area of a triangle with
a base of square root 30 meters and a height of square root 6 meters.

Answer by checkley77(12844)

A=bh/2
A=(sqrt30*sqrt6)/2
A=5.477*2.45/2
A=13.41/2
A=6.7 m. ans.